OpenStudy (anonymous):

A model rocket is launched from the ground with an initial velocity of 352 ft/sec. How long will it take the rocket to reach its maximum height? Show all work in the space provided.

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

use one of the equations for motion at constant acceleration initial velocity u = 352, acceleration a = -32, v = 0 and s = distance so we use v^2 = u^2 + 2as solve for s

OpenStudy (anonymous):

10.9 seconds? @welshfella

OpenStudy (welshfella):

??? - the answer is a distance - there is no time involved in this formula

OpenStudy (anonymous):

It says how long does it take to reach it's maximum height

OpenStudy (welshfella):

oh ! sorry - i misread the question use the v = u + at 0 = 352 - 32t

OpenStudy (welshfella):

t = 352/32 = ?

OpenStudy (anonymous):

11 seconds. @welshfella

OpenStudy (welshfella):

right on

OpenStudy (anonymous):

Thanks :)

OpenStudy (welshfella):

yw